8.

 The charged capacitor in question 7. is then isolated from the battery and a dielectric sheet 

inserted to completely fill the space between the plates. The potential difference between the 

plates drops to 1000 V. Find the dielectric constant of the sheet.


1
Expert's answer
2021-05-21T17:06:05-0400
"C=\frac{\epsilon \epsilon_0 A}{d}\\C_0=\frac{ \epsilon_0 A}{d}\\\epsilon=\frac{C}{C_0}"